# Notes on rational chain connectedness

### v1

@@ 1. Introduction
Let $f\colon X\dashrightarrow Y$ be a meromorphic map between complex analytic spaces, and assume that $(X,\Delta)$ is kawamata log terminal. If $Y$ contains no rational curves, then $f$ is a morphism.
@@
The following corollary is a special case of Theorem 1.6. Consequently, Corollary 1.7 follows from Theorem 1.5 and is independent of Theorem 1.1.
@@ 5. The Hacon–McKernan Rational Chain Connectedness Theorem
We run a minimal model program over $X$ and then apply Lemma 4.4.
@@
In case (B), $(F^{\circ})_{m}$ is contracted by a $(K_{Y_{m}}+(\Gamma^{b}_{\tau^{\circ}})_{m})$-negative divisorial contraction, and hence
$\kappa((F^{\circ})_{m},K_{(F^{\circ})_{m}}+\Phi^{b})=-\infty.$
@@
We may assume $V$ is a common resolution of $F^{\circ}$ and $(F^{\circ})_{m}$:
@@
$p(\operatorname{Nklt}(V,\Delta_{V}))\subset\operatorname{Nklt}(F^{\circ},\Psi),$
so $F^{\circ}$ is rationally chain connected modulo $\operatorname{Nklt}(F^{\circ},\Psi)$, as desired.
@@
Roughly speaking, Hacon and McKernan directly prove that $\kappa_{\sigma}(F^{\circ},K_{F^{\circ}}+\Psi)\leq 0$ by using their difficult extension theorem (see [HM1, Corollary 3.17]). They say that it is the tricky part of their proof of Theorem 1.1 (see [HM2, Theorem 1.2]). In contrast, in our proof of Lemma 5.2, it may happen that
$\kappa(F^{\circ},K_{F^{\circ}}+\Psi)>\kappa\bigl((F^{\circ})_{m},K_{(F^{\circ})_{m}}+\Phi\bigr).$
Thus our method does not allow us to directly deduce $\kappa(F^{\circ},K_{F^{\circ}}+\Psi)\leq 0$. Instead, we prove that a resolution of $((F^{\circ})_{m},\Phi)$ is rationally chain connected modulo the non-klt locus. It then follows that $(F^{\circ},\Psi)$ enjoys the same property.
@@ 6. Proofs of the Results
This theorem is a corollary of Theorem 1.5. In the algebraic setting, the proof is given in [F4, Section 13]. Note that Theorem 1.1 is not needed in this proof. In the complex analytic setting, the same argument applies, since the theory of quasi-log structures is established in [F7]. ∎

### v2

@@ 1. Introduction
Let $f\colon X\dashrightarrow Y$ be a meromorphic map between complex analytic spaces, and assume that $(X,\Delta)$ is divisorial log terminal. If $Y$ contains no rational curves, then $f$ is a morphism.
Corollaries 1.2, 1.3, and 1.4 are stated for divisorial log terminal pairs. However, we note the following simple observation.
Remark 1.5.
Let $(X,\Delta)$ be a kawamata log terminal pair. Then $(X,\Delta)$ is divisorial log terminal.
@@
The following corollary is a special case of Theorem 1.7. Consequently, Corollary 1.8 follows from Theorem 1.6 and is independent of Theorem 1.1 and Theorem 5.1.
@@ 4. A Criterion due to Hacon–McKernan
We also need the following variant of Lemma 4.4 for the proof of Hacon–McKernan’s rational chain connectedness theorem (see Theorem 5.1).
Lemma 4.6.
Let $(X,\Delta)$ be a projective divisorial log terminal pair such that $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier. Let $\varphi\colon X\to W$ be a projective surjective morphism onto a normal projective variety $W$ with connected fibers such that $(K_{X}+\Delta)$ is $\varphi$-ample and $\varphi(\operatorname{Nklt}(X,\Delta))=W$. Let $\psi\colon V\to X$ be a birational morphism from a smooth projective variety $V$ such that both $\operatorname{Exc}(\psi)$ and $\operatorname{Exc}(\psi)\cup\operatorname{Supp}\psi^{-1}_{*}\Delta$ are simple normal crossing divisors on $V$. Write
$K_{V}+\Delta_{V}=\psi^{*}(K_{X}+\Delta)+E,$
where $\Delta_{V}$ and $E$ have no common irreducible components, $\psi_{*}\Delta_{V}=\Delta$, and $E$ is $\psi$-exceptional. Then $V$ is rationally chain connected modulo $\operatorname{Nklt}(V,\Delta_{V})$.
Proof of Lemma 4.6.
Let $H$ be a sufficiently ample Cartier divisor on $W$ such that $(K_{X}+\Delta)+\varphi^{*}H$ is ample. Choose a general effective $\mathbb{Q}$-divisor $A$ on $X$ with sufficiently small coefficients such that
$A\sim_{\mathbb{Q}}-(K_{X}+\Delta)+\varphi^{*}H.$
Set $\Delta_{X}:=\Delta+A$. Then
$K_{X}+\Delta_{X}\sim_{\mathbb{Q}}\varphi^{*}H,$
and hence $\kappa(X,K_{X}+\Delta_{X})\geq 0$. Moreover, since the coefficients of $A$ are sufficiently small and $A$ is general, the pair $(X,\Delta_{X})$ is divisorial log terminal and $A\leq\Delta_{X}$.
Pulling back to $V$, we obtain
$K_{V}+\Delta_{V}+\psi^{-1}_{*}A=\psi^{*}(K_{X}+\Delta_{X})+E.$
Since $A$ is general, we may further assume that
$\operatorname{Exc}(\psi)\cup\operatorname{Supp}\psi^{-1}_{*}A\cup\operatorname{Supp}\psi^{-1}_{*}\Delta$
is a simple normal crossing divisor on $V$.
We now apply Lemma 4.4 to the general fiber of $\varphi\colon X\to W$ with respect to the pair $(X,\Delta_{X})$. The above construction ensures that the assumptions of Lemma 4.4 are satisfied on the general fiber. Therefore we conclude that $V$ is rationally chain connected modulo
$\operatorname{Nklt}(V,\Delta_{V}+\psi^{-1}_{*}A)=\operatorname{Nklt}(V,\Delta_{V}).$
This completes the proof. ∎
@@ 5. The Hacon–McKernan Rational Chain Connectedness Theorem
We run a minimal model program over $X$ and then apply Lemma 4.4 or Lemma 4.6.
@@
Step 3 (Case (A)).
In this step, we treat Case (A).
@@
We may assume that $V$ is a common resolution of $F^{\circ}$ and $(F^{\circ})_{m}$:
[…]
Let $\nu$ be any irreducible component of $\Delta_{V}^{\geq 1}$. Then
$a(\nu,V,\Delta_{V}-E^{\sharp})=a(\nu,V,\Delta_{V})\leq-1.$
Hence, by (5.1), we obtain
$a(\nu,(F^{\circ})_{m},\Phi)\leq-1.$
Since
$K_{(F^{\circ})_{m}}+\Phi=\left(K_{Y_{m}}+(\Gamma_{\tau^{\circ}})_{m}\right)|_{(F^{\circ})_{m}}\sim_{\mathbb{Q}}0,$
it follows that
$a(\nu,F^{\circ},\Psi-(E_{\tau^{\circ}})|_{F^{\circ}})=a(\nu,(F^{\circ})_{m},\Phi)\leq-1.$
Therefore
$a(\nu,F^{\circ},\Psi)\leq-1.$
This shows that
$p(\operatorname{Nklt}(V,\Delta_{V}))=p(\operatorname{Supp}\Delta_{V}^{\geq 1})\subset\operatorname{Nklt}(F^{\circ},\Psi).$
Hence $F^{\circ}$ is rationally chain connected modulo $\operatorname{Nklt}(F^{\circ},\Psi)$.
Step 4 (Case (B)).
In this step, we treat Case (B).
In Case (B), $(F^{\circ})_{m}$ is contracted by a $(K_{Y_{m}}+(\Gamma^{b}_{\tau^{\circ}})_{m})$-negative divisorial contraction. By the $(K_{Y_{m}}+(\Gamma^{b}_{\tau^{\circ}})_{m})$-negative divisorial contraction contracting $(F^{\circ})_{m}$, we obtain a projective surjective morphism
$(F^{\circ})_{m}\to W$
with connected fibers such that $(K_{(F^{\circ})_{m}}+\Phi^{b})$ is ample over $W$ and $\dim W<\dim(F^{\circ})_{m}$. Since
$K_{(F^{\circ})_{m}}+\Phi^{b}\sim_{\mathbb{Q}}(E_{\tau^{\circ}})_{m}|_{(F^{\circ})_{m}}-\Phi^{nb},$
we obtain that $\operatorname{Supp}\Phi^{nb}$ dominates $W$. Hence, the non-klt locus $\operatorname{Nklt}((F^{\circ})_{m},\Phi^{b})$ dominates $W$.
Let $q\colon V\to(F^{\circ})_{m}$ be a resolution such that
$K_{V}+\Delta_{V}=q^{*}(K_{(F^{\circ})_{m}}+\Phi^{b})+E^{\flat},$
We may assume that $V$ is a common resolution of $F^{\circ}$ and $(F^{\circ})_{m}$:
@@
$p(\operatorname{Nklt}(V,\Delta_{V}))\subset\operatorname{Nklt}(F^{\circ},\Psi^{b})=\operatorname{Nklt}(F^{\circ},\Psi).$
Thus $F^{\circ}$ is rationally chain connected modulo $\operatorname{Nklt}(F^{\circ},\Psi)$.
@@
Roughly speaking, Hacon and McKernan directly prove that $\kappa_{\sigma}(F^{\circ},K_{F^{\circ}}+\Psi^{b})\leq 0$ by using their difficult extension theorem (see [HM1, Corollary 3.17]). They say that it is the tricky part of their proof of Theorem 1.1 (see [HM2, Theorem 1.2]). In contrast, in our proof of Lemma 5.2, it may happen that
$\kappa(F^{\circ},K_{F^{\circ}}+\Psi^{b})>\kappa\bigl((F^{\circ})_{m},K_{(F^{\circ})_{m}}+\Phi^{b}\bigr).$
Thus our method does not allow us to directly deduce $\kappa(F^{\circ},K_{F^{\circ}}+\Psi^{b})\leq 0$. Instead, we prove that a resolution of $((F^{\circ})_{m},\Phi)$ (resp. $((F^{\circ})_{m},\Phi^{b})$) is rationally chain connected modulo the non-klt locus. It then follows that the same property holds for $(F^{\circ},\Psi)$ (resp. $(F^{\circ},\Psi^{b})$).
@@ 6. Proofs of the Results
This theorem is a corollary of Theorem 1.6. In the algebraic setting, the proof is given in [F4, Section 13]. Note that Theorem 1.1, and therefore Theorem 5.1, are not needed in this proof. In the complex analytic setting, the same argument applies, since the theory of quasi-log structures is established in [F7]. ∎